OpenStudy (anonymous):

Factor the Following Polynomials: x²-81 x²-13x+36 Which polynomial is a special product and why?

OpenStudy (anonymous):

for the first one use a^2-b^2=(a+b)(a-b) write x^2-81=x^2-(9)^2 now use the above formula

OpenStudy (anonymous):

let me know. what u got ?

OpenStudy (anonymous):

I Don't Get It.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can we write x²-81 as \[x^2-81=(x)^2-(9)^2\]

OpenStudy (anonymous):

reply fast

OpenStudy (anonymous):

Yes You Can Write It as that

OpenStudy (anonymous):

ok now using formula \[a^2-b^2=(a+b)(a-b)\] write \[(x)^2-(9)^2\] as \[(x)^2-(9)^2=(x+9)(x-9)\] is it oky ?

OpenStudy (anonymous):

Yes its okay

OpenStudy (anonymous):

ok now for the second one can i write as \[x^2-9x-4x-36\] ? let me know

OpenStudy (anonymous):

No you have to change the -36 to +36 at the end so would it be x^2-9x-4x+36

OpenStudy (anonymous):

so u are alive:P ok \[x^2-9x-4x+36\] now ok!

OpenStudy (anonymous):

ok take common x from the first two terms and -4 from the last two terms \[x(x-9)-4(x-9)\] are u getting this?

OpenStudy (anonymous):

@kenya.porter

OpenStudy (anonymous):

ok now write above as \[(x-9)(x-4)\] these are the factors hope you got it.

OpenStudy (anonymous):

Oh ok so from (x-9) (x-4) you take out the x's which become x^2 and then -9-4=-13and -9*-4=36 which gives me the equation i started with x^2-13x+36

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank You So Much.

OpenStudy (anonymous):

yw:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

It asked for which polynomial is the special product

OpenStudy (anonymous):

can't you figure out now?

OpenStudy (anonymous):

Would it be x^2-81 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because it is standard formula \[a^2-b^2=(a+b)(a-b)\]

OpenStudy (anonymous):

Yes I Understand it that's why that was my answer. Thank you again so much for the help